The perimeter of a rectangle is 62 m. If the width were doubled and the length were increased by 21 m, the perimeter would be 120 m. What are the length and width of the rectangle?
Answer by richwmiller(17219)

2L+2W=62
2(L+21)+2*(2W)=120
2L+42+4W=120
2L+42+2W+2W=120
2L+2W+2W+42=120
62+42+2W=120
104+2W=120
2W=16
W=8
2L+2*8=62
2L=62-16
divide by 2
L=31-8
L=23
check
2W+2L=62
16+46=62
62=62
ok
check
2*((16)+(23+21))=120
2*(16+44)=120
2*60=120
120=120
ok